# Self-Supervised Alignment with Mutual Information Learning to Follow Principles without Preference Labels

### v1

@@ 6 Experiment 2: Summarization (TL;DR)
6.1 Results: Weak Principle Writer
Post: I decided I couldn’t wait for my ex to come around since there was no guarantee that me waiting for her would be worth it. Sure since the breakup we hadn’t talked as much obviously but now that we are done seemingly forever I can’t comprehend at all knowing that we will never laugh, kiss, talk etc etc together ever again. We had so many plans and so much love but now all that is gone. How do I coupe with that? How can I move on and be happy after knowing all the hard work we put in and all the love is out the door now. I just can’t wrap my head around knowing that the girl I thought I would marry will now move on and live her life away from me till the day we both die […]
Summary mixtral-8x7b: The post is about a person who is struggling to cope with the end of a relationship. They are feeling lost and alone, and are afraid of the future. They are trying to come to terms with the fact that the person they thought they would marry will now move on and live their life away from them. They are still in love with their ex, and are struggling to move on.
Summary mixtral-8x7b + SAMI iteration 3: The post describes the writer’s feelings after a breakup with their ex. They feel lost and alone, and are struggling to move on. They still care for their ex and are afraid of the future.
6.2 Results: Strong Principle Writer
7 Limitations
@@
We proposed a simple method for aligning a pretrained LM with a set of behavioral principles without the need for preference labels or in-context demonstrations. For research purposes, we restricted our experiments to two domains: dialogue and summarization, using a small set of behavioral principles for summarizing Reddit posts or helpful and harmless norms for responding to a wide range of user queries sourced from HH-RLHF. To evaluate the scalability of SAMI to more complex constitutions, future work should include more diverse principles that are representative of personas with diverse preferences [e.g., 9, 22, 12]. Another limitation is that the SAMI loss (Figure 5) requires regularization. Training for too long or failing to regularize can result in forgetting and the model outputting "gibberish", a problem faced by RLHF more generally and usually regularized against using a KL-divergence penalty [e.g., 31]. Moreover, SAMI suffers from a length bias similar to other methods, such as DPO. While our experiments on TL;DR have shown that this length bias can be regularized against by explicitly stating that responses should be concise in the constitution, future extensions could explore incorporating length penalties such as in [26] to avoid a length bias more generally. Furthermore, for SAMI to be effective, the principle-generating model must provide sufficient coverage for contrasts to work, and there must be at least a weak initial connection between the principles and appropriate behavior.
8 Conclusion
@@
SAMI represents progress in teaching a pretrained language model to follow behavioral principles without the use of preference labels, demonstrations, or human oversight. By iteratively finetuning a language model to increase the conditional mutual information between constitutions and self-generated responses given queries from a dataset, SAMI enables the model to connect principles to behavior preferences. Our results demonstrate the potential of this approach: after a small number of gradient updates on self-generated data, the SAMI-trained model outperforms both the initial model and a strong instruction-finetuned baseline on dialogue. On summarization, it surpasses the initial model and performs close to chance when compared to the instruction-finetuned baseline. This success provides evidence that alignment brings out behavioral regularities that the base model implicitly knows, and this insight can be used to craft novel alignment methods.
9 Acknowledgements
This work was supported by a Stanford Human-Centered Artificial Intelligence (HAI) Google Grant and the Microsoft AFMR program. We particularly thank Omar Shaikh and Violet Xiang for helpful and detailed comments, as well as Michael Li and other members of the CoCoLab for useful discussions.
@@ Appendix A Appendix
We provide additional experimental details and prompts. For a reference implementation including further details, see https://github.com/janphilippfranken/sami.
A.1 Hyperparameters
A.1.1 Inference
Across all experiments, we use a temperature of $\tau=0$ when sampling responses from a model. We restrict the maximum sequence length to 350 tokens and use vllm [18] for efficient sampling.
A.2 Training
We use identical settings across all training runs, except for (1) replacing AdamW with RMSprop when finetuning mixtral-8x7b; and (2) using activation checkpointing for finetuning mixtral-8x7b. Importantly, we only use twelve gradient steps overall, starting with two gradient steps at iteration one, followed by four gradient steps at iteration two, and six gradient steps at iteration three. We use FSDP and a custom trainer class for distributed training.
@@ A.2 Training
mistral-7b mixtral-8x7b
# iterations $N$ 3 3
batch size 128 128
$N_{b}$ (# batches at start) 2 2
# gradient steps on each batch 1 1
increment # batches after iteration $N_{b}$ $N_{b}$
gradient steps at each iteration 2, 4, 6 2, 4, 6
# constitutions per query $N_{c}$ 2 2
learning rate $\alpha$ 5e-7 5e-7
precision bf16 bf16
optimizer AdamW RMSprop
@@
# GPUs (A100 80GB) 8 8
sharding strategy Full Shard Full Shard
backward prefetch Backward Pre Backward Pre
auto wrap policy TransformerAutoWrap TransformerAutoWrap
transformer layer class MistralDecoderLayer MixtralDecoderLayer
activation checkpointing No Yes
A.3 PyTorch Implementation
@@
A.4 Uncorrected HH-RLHF Win Rates
@@
A.5 Example Responses HH-RLHF
@@ A.6 Example Responses TL;DR
Post: Me and my girlfriend have been together for about 7 months now, and it has been a blast since the very first day. We basically went from rather good friends to an amazing couple without any awkwardness or friendzone-y (?) problems. Because of this, we are both pretty much convinced that we have found “the one”. However, I recently learned that she is rather envious of my past girlfriends, my last ex in particular. As far as I can tell, her main concern is that I have been intimate in the same ways with them as I have been with her. She can’t quite seem to grasp the idea that with her, everything is very much different. Sexual acts are much more meaningful, if that makes any sense, than they were before. Additionally, she has a hard time processing the idea that I, at some point, loved and cared for them too in the same way as I care for her now. I tend to be a little överly-attachedät times, so I can’t really object to this statement. It is not entirely impossible that I care for her the way I do simply because it’s in my nature. However, I am positive that I have never been in love with anyone as I have been with my current girlfriend, and wouldn’t want it all to end because of this. I have consoled her the best I could, but I’d rather keep it from you redditors. This way, it might be easier to come up with an objective advice.
@@
Post: It is so disheartening to be trying to lose weight and see the scale go up. Sometimes it goes up several days in a row and there you are, first thing in the morning, sleepy and naked and standing on the scale thinking, “What the actual fuck?”. Then you try to shake it off and you start the water running for a shower and you tell yourself, ’It’s okay, it’s just a temporary rise, I will keep trying.’ And then you go on about your day, LIKE THE GODDAMN WEIGHT LOSS HERO THAT YOU ARE. To track my weight loss over the last 4 months I made a graph in a spreadsheet program. It’s not a straight line down, it’s actually peaks and valleys. The *general* trend is down (28lbs today - holla!) but what I’ve noticed is that every time the ẅhoosh fairyv̈isits me and gives me a nice chunk of weight loss, she takes back some of it over the next week. Then she gives me just a *bit* more - enough to keep the graph going down. Thank you, Whoosh Fairy. I love you. I’ve noticed this time and time again so I keep trying, even when the scale goes up despite my best efforts (and even when it goes up after my *worst* efforts).
@@
A.7 Principle Seed Prompt

### v2

@@ 4.2 Experiment 2: Weak and Strong Principle Writer
Figure 3: Experiment 2: Summarization (TL;DR). Conditional MI and Sequence Length. [a] Left: Conditional MI lower bound at each iteration (TL;DR only) for finetuned mistral-7b and mixtral-8x7b for principles written by mistral-7b-instruct. The dashed line indicates the MI for mistral-7b-instruct. Right: Average sequence length for mistral-7b and mixtral-8x7b on the TL;DR dataset using principles written by mistral-7b-instruct. The dashed line represents the sequence length of mistral-7b-instruct. [b] Left: Conditional MI lower bound at each iteration, using the same settings as in [a] but with principles written by claude-opus. Right: Average sequence length, using the same settings as in the right panel of [a], but with principles written by claude-opus. Dashed lines correspond to MI and sequence lengths from the instruct version of a model. Error bars correspond to $\pm$ SEM across 250 data points for all panels.
4.2.1 Results: Weak Principle Writer
Example: Post: I decided I couldn’t wait for my ex to come around since there was no guarantee that me waiting for her would be worth it. Sure since the breakup we hadn’t talked as much obviously but now that we are done seemingly forever I can’t comprehend at all knowing that we will never laugh, kiss, talk etc etc together ever again […] Summary mixtral-8x7b: The post is about a person who is struggling to cope with the end of a relationship. They are feeling lost and alone, and are afraid of the future. They are trying to come to terms with the fact that the person they thought they would marry will now move on and live their life away from them. They are still in love with their ex, and are struggling to move on. Summary mixtral-8x7b + SAMI iteration 3: The post describes the writer’s feelings after a breakup with their ex. They feel lost and alone, and are struggling to move on. They still care for their ex and are afraid of the future.
4.2.2 Results: Strong Principle Writer
@@
4.3 Experiment 3: Scaling to Stronger Models and Diverse Principles
Our previous experiments were limited to a small set of desirable principles, such as helpful and harmless dialogue principles or concise and comprehensive summarization principles, as well as small (mistral-7b) to medium-sized LMs (mixtral-8x7b). To explore whether SAMI generalizes to a more diverse set of principles and larger, more capable models, we finally finetuned llama3-70b on TL;DR by sampling from a diverse set of twenty summarization principles and antitheses (e.g., “talk like a pirate” or “use emojis”; see Section A.13). For this experiment, we approximated Equation 2 by randomly sampling two principles from the list in Section A.13 and then randomly selecting either their definitions or antitheses to generate two contrastive constitutions for each query in the dataset.
Figure 5: Experiment 3: Diverse Summarization Principles. Win rates of the finetuned llama3-70b model against the base model for principles used during training (“train”) and held-out (“test”) principles, with and without chain-of-thought (CoT) (see Section A.16). Error bars correspond to $\pm$ SEM across 250 data points.
During training, we used the first fifteen principles (train), and during evaluation, we further evaluated the final five principles which were held out during training (test). In this experiment, we also explored a chain-of-thought (CoT) variant, which allowed the model to reason about how it would use the principles to summarize before providing a summary (see Section A.16 for the prompt). We expected that more diverse principles would prevent overfitting and allow SAMI to train for longer, so we doubled the number of batches at each iteration compared to our previous experiments (see Section A.3 for detailed hyperparameters). As shown in Figure 5, win rates increased up to 60% ( 68% w/ CoT) for constitutions generated from principles seen during training, and 63% ( 67% w/ CoT) for constitutions generated from held-out principles during the third iteration. These findings suggest that SAMI benefits from stronger models and chain-of-thought reasoning, and that it generalizes well to diverse principles not seen during training. Mutual information and sequence lengths are shown in Figure 8, and example responses are provided in Section A.9.
5 Limitations and Conclusion
@@
Limitations. We restricted our experiments to two domains: dialogue and summarization, using a small set of behavioral principles for summarizing Reddit posts or helpful and harmless norms for responding to a wide range of user queries sourced from HH-RLHF. To further evaluate how well SAMI scales, future work should include more diverse constitutions and tasks, featuring multi-turn interactions, multiple principles, and personas with a wider range of preferences [11, 9, 13, 1]. Training on a broader range of tasks and constitutions is likely to improve a SAMI-trained model’s ability to follow constitutions more consistently and effectively across various domains and scenarios. This could then be evaluated against more capable instruction-following models such as llama3-70b-instruct using benchmarks like MT Bench [41]. A current limitation is that the SAMI loss (Figure 6) requires regularization. Training for too long or failing to regularize can result in forgetting and the model outputting “gibberish”, a problem faced by RLHF more generally and usually regularized against using a KL-divergence penalty [e.g., 30]. Moreover, SAMI suffers from a length bias similar to other methods, such as DPO. While our experiments on TL;DR have shown that this length bias can be regularized against by explicitly stating that responses should be concise, future extensions could explore incorporating length penalties [26]. Furthermore, for SAMI to be effective, the principle-generating model must provide sufficient coverage for contrasts to work, and there must be at least a weak initial connection between the principles and appropriate behavior.
@@
Conclusion. SAMI represents progress in teaching a pretrained language model to follow behavioral principles without the use of preference labels, demonstrations, or human oversight. By iteratively finetuning a language model to increase the conditional mutual information between constitutions and self-generated responses given queries from a dataset, SAMI enables the model to connect principles to behavior preferences. Our results demonstrate the potential of this approach: after a small number of gradient updates on self-generated data, the SAMI-trained model outperforms both the initial model and a strong instruction-finetuned baseline on dialogue (Experiment 1). On summarization, it surpasses the initial model and performs at parity compared to the instruction-finetuned baseline (Experiment 2). Moreover, SAMI benefits from stronger models and more diverse principles, generalizing to principles not seen during training (Experiment 3). This success provides evidence that alignment can leverage the behavioral regularities implicitly learned by the base model.
@@ Appendix A Appendix
A.1 Broader Impacts
We do not foresee any direct societal risks stemming from our work. However, given that the HH-RLHF dataset includes potentially harmful requests and our constitutions included principles that supported the generation of harmful content, we have not uploaded responses to HH-RLHF queries to our online repository to prevent potential misuse. More generally, our approach could support the training of models that can learn to follow harmful principles. Nonetheless, we believe that the ability to follow principles is more likely to benefit the steering of models through system messages that make them more robust to misuse.
A.2 Derivation
We base our objective on the InfoNCE with a tractable conditional from [Eq. 12 in 27]:
$I(X;Y)\geq\mathbb{E}\left[\frac{1}{K}\sum_{i=1}^{K}\log\frac{p(y_{i}|x_{i})}{\frac{1}{K}\sum_{j=1}^{K}p(y_{i}|x_{j})}\right]$ (3)
where the expectation is over $\prod_{i,j}p(x_{i},y_{j})$. Rewriting this to account for the conditional mutual information $I(Y,C;x_{i})$, we get:
$I(Y,C;x_{i})\geq\mathbb{E}\bigg{[}\frac{1}{C}\sum_{j=1}^{C}\log\frac{\pi(y_{ij}|x_{i},c_{j})}{\frac{1}{C}\sum_{k=1}^{C}\pi(y_{ij}|x_{i},c_{k})}\bigg{]}$ (4)
which corresponds to the “contrast over constitutions” (i.e., the second term in Equation 2). We derive our second bound (i.e., “the contrast over responses”) from the symmetry of mutual information:
$I(Y,C;x_{i})\geq\mathbb{E}\bigg{[}\frac{1}{C}\sum_{j=1}^{C}\log\frac{\pi(c_{j}|x_{i},y_{ij})}{\frac{1}{C}\sum_{k=1}^{C}\pi(c_{j}|x_{i},y_{ik})}\bigg{]}$ (5)
by assuming that constitutions are sampled uniformly and independent of queries $x_{i}$:
$\pi(c_{j}|x_{i},y_{ij})=\frac{P(c_{j},y_{ij}|x_{i})}{P(y_{ij}|x_{i})}=\frac{\pi(y_{ij}|x_{i},c_{j})}{P(y_{ij}|x_{i})}P(c_{j}|x_{i})\propto\frac{\pi(y_{ij}|x_{i},c_{j})}{P(y_{ij}|x_{i})}$ (6)
Plugging $\frac{\pi(y_{ij}|x_{i},c_{j})}{P(y_{ij}|x_{i})}$ back into Equation 5 gives us:
$I(Y,C;x_{i})\geq\mathbb{E}\left[\frac{1}{C}\sum_{j=1}^{C}\log\frac{\frac{\pi(y_{ij}|x_{i},c_{j})}{P(y_{ij}|x_{i})}}{\frac{1}{C}\sum_{k=1}^{C}\frac{\pi(y_{ik}|x_{i},c_{j})}{P(y_{ik}|x_{i})}}\right]$ (7)
Assuming that the marginal probability over responses is the same across the sampled responses with constitutions, we can rewrite the above as:
$I(Y,C;x_{i})\geq\mathbb{E}\left[\frac{1}{C}\sum_{j=1}^{C}\log\frac{\pi(y_{ji}|x_{i},c_{j})}{\frac{1}{C}\sum_{k=1}^{C}\pi(y_{ik}|x_{i},c_{j})}\right]$ (8)
which is the same as the “contrast over responses” in Equation 2.
A.3 Hyperparameters
We provide additional experimental details and prompts. For a reference implementation including further details, see our reference implementation using the TL;DR dataset https://github.com/janphilippfranken/sami.
Training. Across all experiments, we use a temperature of $\tau=0$ when sampling responses from a model. We restrict the maximum sequence length to 350 tokens (except for the CoT version of Experiment 3 which includes a longer prompt; here the limit is 500 tokens; see Section A.16). We use vllm [18] for efficient sampling.
Training. In all experiments, we train the initial model $\pi_{\text{BASE}}$ three times on contrastive pairs based on responses sampled from each intermediate model $Q_{1},Q_{2},...,Q_{\eta}$. At each iteration $\eta$, we alternate between two splits of a given dataset to avoid sampling responses to queries present during that model’s training, which could lead to overfitting. We use a batch size of 128 across all experiments and always take one gradient step on each batch. Following previous work [39], we start with a small number of examples at iteration one, specifically 256 for Experiments 1–2 (i.e., num. batches $N_{b}=2$, resulting in two gradient steps, one per batch). We then add two additional batches, resulting in four batches (512 examples, four gradient steps) at iteration two and six batches (768 examples, six gradient steps) at iteration three. In Experiment 3, we train on a more diverse set of constitutions, which is why we double the number of batches at each iteration. For finetuning mistral-7b, we use the AdamW optimizer. For larger models (mistral-8x7b and llama3-70b), we use RMSprop and activation checkpointing. We employ FSDP and a custom trainer class for distributed training (see Table 1 for more details).
@@ A.3 Hyperparameters
mistral-7b mixtral-8x7b llama3-70b
# iterations $N$ 3 3 3
batch size 128 128 128
$N_{b}$ (# batches at start) 2 2 4
# gradient steps per batch 1 1 1
increment # batches $N_{b}$ 2 2 4
# gradient steps per iteration 2, 4, 6 2, 4, 6 4, 8, 12
# constitutions per $N_{c}$ 2 2 2
learning rate $\alpha$ 5e-7 5e-7 5e-7
precision bf16 bf16 bf16
optimizer AdamW RMSprop RMSprop
@@
# GPUs (A100 80GB) 8 8 8
sharding strategy Full Shard Full Shard Full Shard
backward prefetch Backward Pre Backward Pre Backward Pre
auto wrap policy TransformerAutoWrap TransformerAutoWrap TransformerAutoWrap
transformer layer class MistralDecoderLayer MixtralDecoderLayer LlamaDecoderLayer
activation checkpointing No Yes Yes
A.4 PyTorch Implementation
@@
A.5 Uncorrected HH-RLHF Win Rates
@@
A.6 Additional Results for Experiment 3
Mutual information and sequence length for Experiment 3 are shown below.
Figure 8: Experiment 3: MI and Sequence Length. We fine-tune llama3-70b in both panels using principles written with claude-opus. [a] Left: MI lower bound. Right: Average sequence length. Error bars correspond to $\pm$ SEM across 250 data points for both panels. [b] Same as in [a] but now using the CoT prompt during the sampling of responses (see Section A.16).
A.7 Example Responses Experiment 1 (HH-RLHF)
@@ A.8 Example Responses Experiment 2 (TL;DR)
Post: Me and my girlfriend have been together for about 7 months now, and it has been a blast since the very first day. We basically went from rather good friends to an amazing couple without any awkwardness or friendzone-y (?) problems. Because of this, we are both pretty much convinced that we have found “the one”. However, I recently learned that she is rather envious of my past girlfriends, my last ex in particular. As far as I can tell, her main concern is that I have been intimate in the same ways with them as I have been with her. She can’t quite seem to grasp the idea that with her, everything is very much different. Sexual acts are much more meaningful, if that makes any sense, than they were before. Additionally, she has a hard time processing the idea that I, at some point, loved and cared for them too in the same way as I care for her now. I tend to be a little ‘̀overly-attached‘̀ at times, so I can’t really object to this statement. It is not entirely impossible that I care for her the way I do simply because it’s in my nature. However, I am positive that I have never been in love with anyone as I have been with my current girlfriend, and wouldn’t want it all to end because of this. I have consoled her the best I could, but I’d rather keep it from you redditors. This way, it might be easier to come up with an objective advice.
@@
Post: It is so disheartening to be trying to lose weight and see the scale go up. Sometimes it goes up several days in a row and there you are, first thing in the morning, sleepy and naked and standing on the scale thinking, “What the actual fuck?”. Then you try to shake it off and you start the water running for a shower and you tell yourself, ’It’s okay, it’s just a temporary rise, I will keep trying.’ And then you go on about your day, LIKE THE GODDAMN WEIGHT LOSS HERO THAT YOU ARE. To track my weight loss over the last 4 months I made a graph in a spreadsheet program. It’s not a straight line down, it’s actually peaks and valleys. The *general* trend is down (28lbs today - holla!) but what I’ve noticed is that every time the ‘̀whoosh fairy‘̀ visits me and gives me a nice chunk of weight loss, she takes back some of it over the next week. Then she gives me just a *bit* more - enough to keep the graph going down. Thank you, Whoosh Fairy. I love you. I’ve noticed this time and time again so I keep trying, even when the scale goes up despite my best efforts (and even when it goes up after my *worst* efforts).
@@
A.9 Example Responses Experiment 3 (TL;DR) using Diverse Summarization Principles w/ CoT
We report representative responses. Constitution are based on the principles shown in Section A.13.
Summarization Principles:
1. Summaries should use informal, non-technical language.
2. Summaries should rely as heavily as possible on emojis to convey the main ideas and emotions of the original post.
Post: Hey relationship redditors, I have a quick and intriguing question about relationships in college. What would you do in this situation? It is kind of awkward once I really think about it. Here’s my short story. I’m 20 years old in college. In our college’s honor program enjoying life and utilizing my brain to the highest capacity. I’ve met some really cool people over the couple years at my college. I wasn’t too interested in girl’s at first, however I am comfortable around them now. I wasn’t before. I still am a little reserved, but that’s all worked out now. I’ve changed up my lifestyle and negative habits and I’m in college to be a doctor! Only black man in my honors philosophy course and everyone else is white. Representing the black community! We need to represented in these sophisticated classes! XD Anyway, I knew this really nice girl from last semester. We met in the hall after a random instance. I don’t know how we got to know each other, but anyway I really like talking to her. However, she told me that she’s not too interested in doing anything out. Like going out. She’s pretty comfy where she’s at now and doesn’t want to date. She told me that she liked me, but doesn’t want to do that type of thing. It’s pretty understandable. She has a sister and I can sense that she really likes me! 0.o. It is a weird feeling, but all I do is politely bob my head to say hi and she really lights up when she sees me. And I ended up eating lunch with her. It was a really good meal and we have similar interests. She’s a vegetarian and was on a cleanse and ate some really good looking salad! That alone was very admirable! Because I thought, “wow that’s exactly what I’d do!” She kind of hinted and have me good vibes, but I could be wrong."
llama3-70b:
llama3-70b + SAMI Iteration 3:
Summarization Principles:
1. Summaries should be as minimalist as possible, using the fewest words necessary while still preserving the core meaning.
2. Summaries should use only text.
Post: This is my first post so please be kind :) I know that lots of people often feel confused when they come out of a long-term relationship. They think they have forgotten how to be single, or how to flirt/date. I am one of these people. The problem is, my relationship started when I had just turned 16. I have never been single - as an adult. That might sound silly. But the only time I have ever flirted or dated was as an over-confident, hormone-riddled teenager. Now I have a pretty demanding job, responsibilities blah blah… And I just don’t know how to this! I’m no way in a rush to get into a new relationship, but that doesn’t mean I want to be completely alone in the mean time. If anyone has experienced anything similar, or just generally has some advice, it would be greatly appreciated!"
llama3-70b: The post is about a person who is struggling to adjust to being single after a long-term relationship. They are looking for advice on how to flirt and date as an adult.
llama3-70b + SAMI Iteration 3: Person feels confused after long-term relationship. Wants advice on how to flirt/date.
Summarization Principles:
1. Summaries should be as empathetic as possible, demonstrating a deep understanding of the original poster’s perspective, feelings, and intentions.
2. Summaries should incorporate as much distinctive pirate jargon, phrases, and style as possible to add a creative, engaging twist.
Post: We broke up a few days before christmas and I have been very upset and sad since then. Basically immediately after the breakup he started seeing someone new which really hurt me of course cause I had hope we would get back together. Since then I have tried seeing other people. Things always seem to be going well at first but then I feel like the other person doesn’t really care that much or want to date me. I feel extremely lonely and miss being in a relationship. I really miss the companionship, and I know I should become comfortable with being on my own but it really sucks and I feel miserable. Especially when I think about my ex and how he is doing everything we used to do together with someone else. Then this just makes me feel worse cause no one seems to want to do anything with me. I feel like people and especially my ex are almost laughing at me that I haven’t found anyone new. And it seriously seems like everyone else is in a happy relationship while I am sitting here preparing to be alone forever. I know I shouldn’t be trying to find happiness in a relationship, I just don’t want to end up alone forever.",
llama3-70b: Ahoy there, matey! It sounds like ye’ve been through a rough patch, what with the breakup and all. It’s no wonder ye feel lonely and miss the companionship of a relationship. But don’t ye worry, there be plenty of fish in the sea! And as for yer ex, he be a scurvy dog for movin’ on so quickly. But don’t let that get ye down, matey. Ye be a fine catch, and I’m sure ye’ll find someone who’ll appreciate ye for who ye are. In the meantime, try to enjoy yer freedom and focus on yerself. Ye don’t need a relationship to be happy, and ye certainly don’t want to end up with someone who doesn’t truly care for ye. So keep yer head up, matey, and remember that ye be a strong, independent pirate who don’t need no man to make ye happy!
llama3-70b: + SAMI Iteration 3: Ahoy matey, ye be feelin’ like a lonely scallywag since ye broke up with yer ex. Shiver me timbers, it be hurtin’ to see ’em walkin’ the plank with someone new. Ye be tryin’ to find a new crew, but it seems like no one be wantin’ to join yer ship. Ye be missin’ the companionship, and it be makin’ ye feel miserable. Especially when ye think about yer ex doin’ all the things ye used to do together with someone else. It be makin’ ye feel like everyone be laughin’ at ye, and ye be feelin’ like ye’ll be alone forever. But don’t give up hope, matey, ye don’t need a relationship to find happiness, ye just don’t want to end up walkin’ the plank alone.
A.10 Principle Seed Prompt